HIGHER VIBRATIONAL LEVELS OF THE Ã^{1}Π_{u} STATE OF C_{3} OBSERVED BY LASER-INDUCED FLU-ORESCENCE

CHIAO-WEI CHEN, \underline{ANTHONY J. MERER}, JUN-MEI CHAO, AND YEN-CHU HSU, Institute of Atomic and Molecular Sciences, Academia Sinica, Box 23-166, Taipei, Taiwan 10617.

The vibrational structure of the Ã^{1}Π_{u} electronic state of C_{3} in the region 26000-31000 cm^{−1} has been re-examined, using laser excitation spectra of jet-cooled molecules. Rotational constants and vibrational energies have been determined for over 60 previously unreported vibronic levels; a number of other levels have been re-assigned. The vibrational structure is complicated by interactions between levels of the upper and lower Born-Oppenheimer components of the Ã^{1}Π_{u} state, and by the effects of the double minimum potential in the Q_{3} coordinate, recognized by Izuha and Yamanouchi. ^{a} The present work shows that there is also strong anharmonic resonance between the overtones of the ν_{1} and ν_{3} vibrations. For instance, the Σ_{u}^{+} vibronic levels 2 1^{+}1 and 0 1^{+}3 are nearly degenerate in zero order, but as a result of the resonance they give rise to two levels 139 cm^{−1} apart, centered about the expected position of the 2 1^{+}1 level. Similarly, the 202 level lies 60 cm^{−1} lower than expected because of interaction with the 400 and 004 levels. With these irregularities recognized, every observed vibrational level up to 29550 cm^{−1} (a vibrational energy of nearly 5000 cm^{−1}) can now be assigned.

aM. Izuha and K. Yamanouchi, J. Chem. Phys. 109, 1818 (1998).